OpenStudy (anonymous):

PQ = |9 - (-3)| = 12 QR = |-2 - (7)| = -9 PR = sqrt(9 - (-3))^2 + (-2 - (7))^2 = sqrt 12^2 + (-9)^2 = sqrt 144 + 81 = sqrt 225 = 15 12 + 15 + (-9) = 18 The perimeter of PQR is 18 units. Does this problem look correct? Ruler postulates are fun. :D

OpenStudy (anonymous):

\[PQ = |9 - (-3)| = 12\]\[QR = |-2 - (7)| = -9\]\[PR = \sqrt(9 - (-3))^2 + (-2 - (7))^2\]\[[\sqrt (12^2) + (-9)^2]\]\[(\sqrt 144 + 81)\]\[\sqrt 225\]\[= 15\]\[12 + 15 + (-9) = 18\]

OpenStudy (anonymous):

Something looks wrong here. Does anyone know?